radomes.openskiesproject.org
hghiafkjdg hghiafkjdg
  • 09-06-2016
  • Mathematics
Answered

A coin is weighted so that the probability of getting heads is 0.70. If the coin is flipped twice, what is the probability of getting a tail followed by a head?

Verified Answer :

caylus
caylus caylus
  • 09-06-2016

Hello,

p=3/10*7/10=21/100=0.21
====================

Answer Link

Related Questions

A box of chalk is in the shape of a rectangular prism with a length of 16 centimeters a width of 512 centimeters and a height of 3 centimeters What is the volum
What was one difference between the encomienda system and the repartimiento system of Spanish colonial America A The encomienda system gave Spanish colonists co
Last Summer the lifeguards rescued 6 adult swimmersdirect object
Which character is described in these lines 10 points Then the mighty warspirit endured for a season Bore it bitterly he who bided in darkness That lighthearted
How did the death of Malik Shah affect the Seljuk Empire
how do you find weight massgravitational acceleration
Why have the pyramids of Egypt survived whereas the ziggurats of Mesopotamia have not
fc 95c 32 I just dont understand how to find the inverse of the function
Why do you think AES is a better encryption algorithm
Name the term in soccer that is used when a player stops a pass hisher feet